OpenStudy (sooobored):

QUICK CONVERSION 1000 F -> Celsius

OpenStudy (sooobored):

yes i know i can google it, but dont feel like it

OpenStudy (sooobored):

\[\frac{9}{5} C+32=F\] or other way around

OpenStudy (rcp031):

0 degree Celsius is 32 degree Fahrenheit.

OpenStudy (sooobored):

yes, now give me 1000 fahrenheit in celcius

OpenStudy (rcp031):

(5/9)*1000-32=523.55 C

OpenStudy (sooobored):

eh, i got 537 but close enough